OpenStudy (anonymous):

To solve the equation r divided by four minus two-ninths equals four-sixths for r, what two operations must be performed, and in which order? Add then multiply. Add then divide. Subtract then multiply. Subtract then divide.

OpenStudy (anonymous):

subtract then multiply

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

you have to switch it around it should be add then divide

OpenStudy (anonymous):

because you should all ways turn it around then its easier

OpenStudy (math&ing001):

Actually @iceicebaby you substract first to get the 'r' in one side then you multiply.

OpenStudy (anonymous):

Correct so I was right

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

It is subtract then multiply

OpenStudy (anonymous):

that is wrong because i put that on my test and got it wrong

OpenStudy (math&ing001):

r/4 - 2/9 = 4/6 r/4 - 2/9 + 2/9 = 4/6 + 2/9 r/4 = 8/9 r/4 * 4 = 8/9 * 4 r = 32/9 They expect you to put add then multiply (option 1) But really those notions are just the same because multiplying by 1/4 is equivalent to dividing by 4.